Hello! I am a 7th grade pre-algebra student who is very confused.There is a word problem i have been trying to solve all night. Here it goes!
In the parking lot at a truck stopthere were 6 more cars than 18-wheel trucks. There were 134 wheels in the parking lot.How many cars and trucks were there?
I would greatly appreciate your help.Thank you!
~Morgan Philley~

Answer by stanbon(75887)

In the parking lot at a truck stop there were 6 more cars than 18-wheel trucks. There were 134 wheels in the parking lot.How many cars and trucks were there?
I would greatly appreciate your help.Thank you!
---------------
Let the number of trucks be "x".
These x trucks have 18x wheels.
--------------------
Then the number of cars is "x+6"
These x+6 cars have 4(x+6) = 4x+24 wheels.
--------------
EQUATION:
truck wheel + car wheels = 134 wheels
18x + 4x+24 =134
22x+24 = 134
22x = 110
x=5 (there are 5 trucks)
x+6=11 (there are 11 cars)
-----------
Let's see if this is correct.
five 18-wheelers would have 5*18= 90 wheels
eleven cars would have 11*4=44 wheels
And 90 + 44 = 134 wheels.
--------------
Cheers,
Stan H.